A magnesium ribbon weighing 1.25g formed 3.45 of oxide when made to react with oxygen.&
What is the % of oxygen in the compound?
1
Expert's answer
2011-09-30T10:36:13-0400
m(Mg) = 1.25g
m(oxide) = 3.45g
m(O) = m(oxide) -m(Mg) = 3.45 - 1.25 = 2.2 g
w(O) = m(O)/m(oxide)*100% = 2.2/3.45*100% = 63.77%